A 544-kg satellite is in a circular orbit about Earth at a height above Earth equal to Earths mean radius. (a) Find the satellites orbital speed. m/s (b) Find the period of its revolution. (c) Find the gravitational force acting on itA satellite of Mars, called Phobos, has an orbital radius of 9.4 x 106 m and a period of 2.8 104 s. Assuming the orbit is circular, determine the mass of Mars. x 10 s. Assuming kg
Answer #1

F = mv^2/r = GMm/r^2